# Random Schreier graphs as expanders

### v6

@@ A.2 Proof of Theorem 2
Proof.
We proceed in a similar way as in the proof of lemma A.1.4. Consider we have $i$ pairs of matrices that appear exactly twice in $\omega$. In the bipartite setting, the signs are forced by the parity of the position of each letter. We thus choose to ignore them. Then, the probability of having no pair $(s,t)$ such that we have the pattern $\omega=\dots s\dots t\dots s\dots t\dots$ is $\frac{2^{i}}{(i+1)!}$. The proof is the same as that of claim 2 in Lemma A.1.4, except that the numerator and the denominator of the fraction are both divided by $2^{i}$ (because we ignore the signs).
Using the proof of Lemma A.1.2, we conclude that, if a letter appears twice with the same sign, the probability of having a closed walk is less than $\frac{2}{n}$. If there is a pair $(s,t),\omega=\dots s\dots t\dots s^{-1}\dots t^{-1}\dots$, then, by using the argument from lemma A.1.4 (claim 1), the probability of getting a closed walk is $\frac{1}{n}$. If those cases do not happen, we still can bound the probability of getting a closed walk using lemma A.1.3 when at least a letter appears three times. The conditional probability of getting a closed walk is then less than $\frac{5}{n}$.
@@
  • The size of the event in which this bound does not apply, but we can apply the bound from lemma A.1.3 (which is $\frac{5}{n}$) can be computed as follows. $(2i)!\binom{2m}{2i}$ is the number of ways of placing $i$ pairs of letters with different sign in $2m$ positions. Since we ignore the sign, this quantity has to be divided by $2^{i}$ which gives $(\frac{1}{2})^{i}\frac{(2m!)}{(2m-2i)!}$. A fraction $\frac{2^{i}}{(i+1)!}$ of them are well formed. $y_{3}(i)$ is the numbers of ways of filling the remaining gaps so that no letter appear once nor twice, and at least letter appears three times. Choosing the $i$ pairs among the $d$ possible ones and summing over all $i\leq 2m$, we get
@@
Summing all the above probabilities and substituting this in 7 finishes the proof of Theorem 2.
A.3 Proof of Proposition 1
@@
is also an eigenvector with associated eigenvalue $-\lambda$. Thus, $v-v^{\prime}$ is an eigenvector of $P^{2}$ of eigenvalue $\lambda^{2}$ and this vector has $n_{2}$ zeros on the right. Because $P^{2}$ represents a disconected graph, if we reduce the dimension of this vector by $n_{2}$ (removing the zeros on the right corresponding to one connected component) we get an eigenvector of $M.^{t}M$ of eigenvalue $\lambda^{2}$. Therefore, $M.^{t}M$ has the same eigenvalues —denoted $\mu_{1}\geq\mu_{2}...\geq\mu_{n_{1}}$— as $P$, but squared. ∎
The proof works the same with $Q$ (taking $n_{1}=n_{2}$). We note $\alpha$ the bound for $|\mu_{3}(Q)|$ proven in the preceding section. $\alpha$ might refer to the bound from Theorem 1 if the graph is obtained from a bipartite regular graph of even degree or to the bound from Theorem 2 if its degree is odd. In ${}^{t}A.A$, the second largest magnitude eigenvalue is thus $|\lambda_{2}(^{t}A.A)|=(d_{1}|\mu_{3}(Q)|)^{2}\leq d_{1}^{2}\alpha^{2}$. Hence we have
@@
A.4 Remarks on the proofs techniques
In section A.1, we have seen that the probability for the random walk to collapse to the identity sequence is less than $(\frac{2}{d})^{m}$. If so, the probability of getting a closed walk (conditioned by the collapsing event) is then 1. When the collapse does not happen, we cannot hope for a smaller probability than $\frac{1}{n}$ to get a closed path in the graph. This is the smallest probability of a closed walk one can get in a random graph. The trace method gives then
@@ A.4 Remarks on the proofs techniques
With $m=\Omega(\ln n)$, we get $\mathbf{E}(|\mu_{2}|)=\mathcal{O}(d^{-\frac{1}{2}})$, which is bigger than the bound from [11] only by a constant factor. This suggests that this technique can be improved by a subtler subdivision of the probability space of $\omega$, as well as a more careful analysis of the probability of having a closed walk (specially with the condition $X_{2}^{\prime}\cap X_{4}$ for $X_{2}^{\prime}$ and $X_{4}$ defined on page A.1).
Let us observe that the term $n^{\frac{1}{2m}}$ is getting close to $1$ only when $m=\Omega(\log n)$. This is why in [6] or [3], the length of the random walk ($2m$) is logarithmic in the number of vertices. Our computations show that the optimal size of the walk should be a bit smaller; this might be because it allows us to assume that, with the overwhelming probability, at least one letter appears in $\omega$ exactly one time (the event $X_{1}$ in the proof of Theorems 1 and 3). Clearly, we cannot keep $m$ small and at the same time make the factor $n^{\frac{1}{2m}}$ close to $1$. This is an important limitation of our technique.
@@
Conjecture 1.
Let $G_{k}=Sch(GL_{k}(\mathbb{F}_{q})\circlearrowleft(\mathbb{F}_{q}^{k})^{*},S)$ with $S$ a random subset of $GL_{k}(\mathbb{Z}_{q})$ and $q$ a prime number. Let $G^{\prime}_{k}$ be a $2|S|$-regular graph from the permutation model of size $q^{k}-1$. Then, as $k$ grows, the second largest eigenvalue distribution of $G$ converges to that of $G^{\prime}$.

### v7

@@ 0.B.1 Proof of Proposition 1
Proof
We proceed in a similar way as in the proof of lemma 5. Consider we have $i$ pairs of matrices that appear exactly twice in $\omega$. In the bipartite setting, the signs are forced by the parity of the position of each letter. We thus choose to ignore them. Then, the probability of having no pair $(s,t)$ such that we have the pattern $\omega=\dots s\dots t\dots s\dots t\dots$ is $\frac{2^{i}}{(i+1)!}$. The proof is the same as that of claim 2 in Lemma 5, except that the numerator and the denominator of the fraction are both divided by $2^{i}$ (because we ignore the signs).
Using the proof of Lemma 3, we conclude that, if a letter appears twice with the same sign, the probability of having a closed walk is less than $\frac{2}{n}$. If there is a pair $(s,t),\omega=\dots s\dots t\dots s^{-1}\dots t^{-1}\dots$, then, by using the argument from lemma 5 (claim 1), the probability of getting a closed walk is $\frac{1}{n}$. If those cases do not happen, we still can bound the probability of getting a closed walk using lemma 4 when at least a letter appears three times. The conditional probability of getting a closed walk is then less than $\frac{5}{n}$.
@@
  • The size of the event in which this bound does not apply, but we can apply the bound from lemma 4 (which is $\frac{5}{n}$) can be computed as follows. $(2i)!\binom{2m}{2i}$ is the number of ways of placing $i$ pairs of letters with different sign in $2m$ positions. Since we ignore the sign, this quantity has to be divided by $2^{i}$ which gives $(\frac{1}{2})^{i}\frac{(2m!)}{(2m-2i)!}$. A fraction $\frac{2^{i}}{(i+1)!}$ of them are well formed. $y_{3}(i)$ is the numbers of ways of filling the remaining gaps so that no letter appear once nor twice, and at least letter appears three times. Choosing the $i$ pairs among the $d$ possible ones and summing over all $i\leq 2m$, we get
@@
Summing all the above probabilities and substituting this in 7 finishes the proof of Theorem 0.A.2.
0.B.4 Proof of Proposition 2
@@
is also an eigenvector with associated eigenvalue $-\lambda$. Thus, $v-v^{\prime}$ is an eigenvector of $P^{2}$ of eigenvalue $\lambda^{2}$ and this vector has $n_{2}$ zeros on the right. Because $P^{2}$ represents a disconected graph, if we reduce the dimension of this vector by $n_{2}$ (removing the zeros on the right corresponding to one connected component) we get an eigenvector of $M.^{t}M$ of eigenvalue $\lambda^{2}$. Therefore, $M.^{t}M$ has the same eigenvalues —denoted $\mu_{1}\geq\mu_{2}...\geq\mu_{n_{1}}$— as $P$, but squared.
The proof works the same with $Q$ (taking $n_{1}=n_{2}$). We note $\alpha$ the bound for $|\mu_{3}(Q)|$ proven in the preceding section. $\alpha$ might refer to the bound from Theorem 0.A.1 if the graph is obtained from a bipartite regular graph of even degree or to the bound from Theorem 0.A.2 if its degree is odd. In ${}^{t}A.A$, the second largest magnitude eigenvalue is thus $|\lambda_{2}(^{t}A.A)|=(d_{1}|\mu_{3}(Q)|)^{2}\leq d_{1}^{2}\alpha^{2}$. Hence we have
@@
0.B.5 Remarks on the proofs techniques
In section 0.B.2, we have seen that the probability for the random walk to collapse to the identity sequence is less than $(\frac{2}{d})^{m}$. If so, the probability of getting a closed walk (conditioned by the collapsing event) is then 1. When the collapse does not happen, we cannot hope for a smaller probability than $\frac{1}{n}$ to get a closed path in the graph. This is the smallest probability of a closed walk one can get in a random graph. The trace method gives then
@@ 0.B.5 Remarks on the proofs techniques
With $m=\Omega(\ln n)$, we get $\mathbf{E}(|\mu_{2}|)=\mathcal{O}(d^{-\frac{1}{2}})$, which is bigger than the bound from [12] only by a constant factor. This suggests that this technique can be improved by a subtler subdivision of the probability space of $\omega$, as well as a more careful analysis of the probability of having a closed walk (specially with the condition $X_{2}^{\prime}\cap X_{4}$ for $X_{2}^{\prime}$ and $X_{4}$ defined on page 0.B.2).
Let us observe that the term $n^{\frac{1}{2m}}$ is getting close to $1$ only when $m=\Omega(\log n)$. This is why in [6] or [3], the length of the random walk ($2m$) is logarithmic in the number of vertices. Our computations show that the optimal size of the walk should be a bit smaller; this might be because it allows us to assume that, with the overwhelming probability, at least one letter appears in $\omega$ exactly one time (the event $X_{1}$ in the proof of Theorems 0.A.1 and 1). Clearly, we cannot keep $m$ small and at the same time make the factor $n^{\frac{1}{2m}}$ close to $1$. This is an important limitation of our technique.
@@
Conjecture 1
Let $G_{k}$ be the Schreier graph of $S\subset\textit{GL}_{k}(\mathbb{F}_{q})$ acting on$(\mathbb{F}_{q}^{k})^{*}$ with $S$ a random subset of $\textit{GL}_{k}(\mathbb{Z}_{q})$ and $q$ a prime number. Let $G^{\prime}_{k}$ be a $2|S|$-regular graph from the permutation model of size $q^{k}-1$. Then, as $k$ grows, the second largest eigenvalue distribution of $G$ converges to that of $G^{\prime}$.
@@
Acknowledgements.
The author thanks Andrei Romashchenko and Alexander Shen for providing several scientific ideas and help in the writing process.